# Training Machine Learning models at the Edge: A Survey

### v2

@@ IX-D Lack of labelled data availability
In the context of training ML models at the edge, a prominent challenge arises from the inherent abundance of unlabeled data gathered by edge devices. This issue becomes particularly pertinent as the majority of ML applications traditionally emphasize supervised learning paradigms, necessitating labeled datasets for effective model training and adaptation on edge devices [32]. To address this challenge and pave the way for future research, it is imperative to explore innovative directions. This includes delving into unsupervised (section VI-A), self-supervised (section VI-D), or semi-supervised (section VI-C) techniques and exploring methodologies that can make the most of limited labeled instances such as few-shot learning. Moreover, the development of auto-labeling systems, exemplified by solutions like Flame [367], emerges as a promising avenue to autonomously generate labeled data, mitigating the impact of the labeled data scarcity at the edge.
IX-E Reducing Energy consumption and Carbon Footprint
The surge in distributed learning methodologies, while propelling advancements in ML, has concurrently raised concerns regarding their environmental impact. The substantial energy requirements for training models and data transfer to/from centralized data centers contribute to a significant carbon footprint [368, 369]. Recent trends in machine learning underscore the critical need to estimate and minimize the environmental impact model training processes. To contextualize the magnitude of this issue, according to [370], the estimated carbon footprint associated with edge devices by 2027 will be between 22 and 562 MtCO2-eq/year. Therefore, as a pressing research direction, there is a growing emphasis on developing techniques for training ML models at the edge with meticulous attention to energy efficiency and carbon footprint reduction. Pioneering works, such as [371], have initiated analyses to quantify the environmental impact of ML in edge devices. Notably, frameworks like FedERA [362], designed for training FL models at the edge, incorporate a dedicated carbon emission tracker module to precisely estimate CO2 emissions during the local parameter update phase. More research and development into this, ensures that advancements in Edge ML align with sustainability goals, making it an essential facet of ongoing research in the field.
IX-F Security concerns for distributed learning methods
The advent of distributed learning methodologies, while transformative, introduces notable security concerns, particularly regarding the vulnerability of edge devices [372, 373] and the inherent challenges in securing IoT and mobile devices further compound these concerns [374]. As the cybersecurity landscape for ML models continues to evolve, with a growing literature on the subject (e.g., see [375]), there is a pressing need for comprehensive research into the security implications of training ML models at the edge. Recent studies, exemplified by works like [376, 377, 372, 373], underscore the urgency of understanding potential security vulnerabilities in EL methodologies. Addressing these concerns necessitates further exploration of security measures to be integrated into the training processes of machine learning at the edge, ensuring a robust and secure deployment of distributed learning methods.
IX-G Training Large models at the edge
@@
Large models, including Large Language Models (LLMs) [378, 379, 380], Diffusion models [381], and Audio Generation models [382, 383], etc, are increasingly prevalent and are steadily growing in popularity. However, despite a growing demand for personalization of these models [384, 385] and privacy concerns in collecting and using personal or private data on centralized cloud servers [386]. The fine-tuning, training or personalization of such models at the edge is very challenging considering the limited resources available for edge devices [387]. Although contributions in this domain are currently limited, the predicted surge in interest prompts a need for proactive exploration. Notable approaches, such as FedLLM  [^5] FedLLM is a platform to Build Large Language Models on Proprietary Data using federated learning using the FedML Platform https://doc.fedml.ai/federate/fedllm, FwdLLM [81] and FATE-LLM [80], have emerged using FL to address the challenges of training LLMs at the edge. Looking ahead, the increasing popularity of LLMs and diffusion models anticipates a growing interest in adapting them for EL. Furthermore, innovative techniques such as LoRa [388] and Fnet [389] offer potential solutions for the resource constraints on edge devices, especially when integrated with complementary approaches like FL V-A1, SL V-A2, or model compression techniques V-C. The convergence of these methodologies holds promise for overcoming challenges associated with training large models at the edge in the foreseeable future.
IX-H Extension to privacy preserving application
@@
The escalating popularity of ML applications has brought forth heightened concerns regarding the vast amounts of private and personal data required for effective model training, raising questions about various legal and ethical implications. As discussed in earlier sections, edge learning emerges as a potential solution to address these privacy concerns, as it enables the training of ML models directly on the edge device, eliminating the need for sensitive data to traverse external networks. As such, the usage of edge learning for privacy-preserving applications is expected to be a pivotal research direction for the field. Domains like healthcare (VII-A) often had legal requirements as well as ethical concerns of using the data for training ML model [390]. And as explored in discussed in previous sections (VII-D), recommendation systems also stand out as a promising avenue for exploration because the significant scrutiny faced for their reliance on private data during model training [312, 311]. Additionally, other applications that require model personalization and tuning on private data ranging from spam detection in SMS and emails to word suggestions in keyboards and personal assistant chatbots or HAR, can benefit from edge learning, fostering a paradigm shift towards more ethical and privacy-conscious ML applications.
IX-I Frameworks to implement training
@@
Despite the proliferation of frameworks and libraries aimed at enabling ML on edge devices, the current landscape lacks robust support for on-device training. Existing tools, such as PyTorch Mobile [351] and ExecuTorch [352], predominantly emphasize inference on mobile/edge devices, neglecting the essential backpropagation algorithms crucial for the training phase. This imbalance in focus between training and inference highlights a critical gap in the current ecosystem. Although some tools have been proposed to facilitate on-device training at the edge VIII-A, there is a pressing need for the development of new libraries, frameworks, and tools explicitly designed for edge learning. Foreseeing the surge in interest, we anticipate an influx of innovative solutions in the coming years, addressing the scarcity of options for on-device ML training. Furthermore, this surge in interest is likely to catalyse advancements in IoT and mobile hardware, optimizing their design and capabilities for ML training on the edge.
X Conclusion
In this survey, we tried to cover the extensive field of Edge learning (EL), which we define as the training and tuning of ML models in the edge, we made an attempt to define the metrics and requirements of EL, and explored then compared the different techniques and methodologies used to optimize ML training at the edge. Next we explored the usage and integrations of ML types such as unsupervised learning, reinforcement learning, etc, and the different applications and use cases of EL. We then explored the tools, libraries and frameworks used for simulating EL and training ML models in edge devices. Finally we identified key challenges in EL and attempted to predict the future trends and research directions of the field. For the techniques and methodologies used to optimize EL, we identified that distributed learning methods such as Federated Learning (FL) are more popular and still growing in popularity for EL, which comes as no surprise considering the potential of these techniques in terms of training large ML models in resource constrained devices. We also estimated the benefits and drawbacks of most of the techniques presented in the survey and identified the growing trend of combining different techniques to alleviate their negatives and make the most out of their benefits. However, the scope of this survey is on the broad concept of EL rather than providing an in depth evaluation and comparison of the performances on specific tasks and problems, as such, more research on specific domains, tasks and techniques is required in the future to complement this survey, and this paper can be used as a reference for developing a broad understanding of edge learning, its requirements, challenges, uses cases and trends, as well as, identifying and gaining a broad overview of the principal techniques used to optimize ML models for training at the edge and pinpointing the different tools, simulations, libraries and frameworks for the training in edge devices.

### v3

@@ VII-A3 Lack of labelled data availability
In the context of edge learning, a prominent challenge arises from the prevalence of unlabeled data on edge devices. This issue becomes particularly problematic as the majority of ML applications traditionally emphasize supervised learning paradigms, necessitating labeled datasets for effective training [45]. To address this challenge, it is imperative to explore no-label or few-label solutions. This includes a focus on unsupervised (Section IV-A), self-supervised (Section IV-D), or semi-supervised (Section IV-C) techniques as well as methods that can make the most of limited labeled instances such as few-shot learning. Moreover, the development of auto-labeling systems, exemplified by solutions like Flame [391], emerges as a promising solution to mitigate the impact of the labeled data scarcity on the edge.
VII-A4 Abundance of non-iid data on the edge
As described in Section III-A, distributed learning methods, such as FL, stand out as pivotal techniques in edge learning. However, a significant body of literature on FL is done under the assumption of IID data [392], and very often this assumption doesn’t reflect the data present on edge devices. While the effects of Non-IID data on FL depends widely on the type of FL method employed, it often negatively impact the training [393, 394, 395]. And while multiple solutions have been proposed to handle Non-IID data, they might come at the expense of privacy preservation and a clear benchmark of real Non-IID performance is unclear considering the large diversity in FL methods [393]. More details on the impact of Non-IID data on FL as well as the specific challenges it poses, and the different approaches proposed to tackle the issue can be found on the following surveys [392, 393].
VII-A5 Data leakage and privacy concerns
Although edge learning aims to provide a privacy-aware alternative to traditional machine learning, it is not immune to privacy and leakage risks during or as a result of the training phase. Notably, model vulnerabilities can be exploited to leak sensitive data, particularly in FL settings where the model is shared among clients [396, 397]. This vulnerability is exacerbated when the model is susceptible to data reconstruction attacks, increasing the risk of private data leakage [396]. To mitigate these risks, techniques like differential privacy [101, 100], encryption methods [398], and other approaches are often employed to ensure optimal privacy. However, further research is necessary to guarantee the preservation of private data and minimize leakage risks in edge learning.
VII-B Future trends and research directions
VII-B1 Hybridization of ML techniques in edge learning
In the exploration of edge learning techniques detailed in Section III, various strategies have been employed to optimize the training of relatively large ML models on edge devices. However, each of these techniques, as outlined in Table II, comes with distinct advantages and drawbacks. Recognizing the diversity in these methodologies, there has been a notable surge in approaches that advocate for a hybridization of multiple techniques. Figure 4 illustrates this promising trend, wherein researchers aim to maximize the advantages and mitigate the drawbacks of individual techniques by combining them. This emerging direction, signifies a deliberate effort to create comprehensive and robust solutions tailored to the unique challenges posed by edge devices. Given the promising results showcased by such hybrid approaches, the trajectory indicates a continued surge in interest and research efforts towards refining and expanding the applicability of hybridized techniques in the domain of edge learning.
VII-B2 Training large models at the edge
@@
Large models, including Large Language Models (LLMs) [399, 400, 401], Diffusion models [402], and Audio Generation models [403, 404], etc., are increasingly prevalent and are steadily growing in popularity. However, despite a growing demand for personalization of these models [405, 406] and privacy concerns in collecting and using personal or private data on centralized cloud servers [407]. The fine-tuning, training or personalization of such models at the edge is very challenging considering the limited resources available for edge devices [408]. Although contributions in this domain are currently limited, the predicted surge in interest prompts a need for proactive exploration. Notable approaches, such as FedLLM [^13] FedLLM is a platform to Build Large Language Models on Proprietary Data using FL using the FedML Platform https://doc.fedml.ai/federate/fedllm, FwdLLM [85] and FATE-LLM [84], have emerged using FL to address the challenges of training LLMs at the edge. Looking ahead, the increasing popularity of LLMs and diffusion models anticipates a growing interest in adapting them for edge learning. Furthermore, innovative techniques such as LoRa [409] and Fnet [410] offer potential solutions for the resource constraints on edge devices, especially when integrated with complementary approaches like FL III-A1, split learning III-A2, or model compression techniques III-C. The convergence of these methodologies holds promise for overcoming challenges associated with training large models at the edge in the foreseeable future.
VII-B3 Extension to privacy preserving applications
@@
The escalating popularity of ML applications has brought forth heightened concerns regarding the vast amounts of private and personal data required for effective model training, raising questions about various legal and ethical implications. As discussed in earlier sections, edge learning emerges as a potential solution to address these privacy concerns, as it enables the training of ML models directly on the edge device, eliminating the need for sensitive data to traverse external networks. As such, the usage of edge learning for privacy-preserving applications is expected to be a pivotal research direction for the field. Domains like healthcare (V-A) often had legal requirements as well as ethical concerns of using the data for training ML model [411]. And as explored in discussed in previous sections (V-D), recommendation systems also stand out as a promising avenue for exploration because the significant scrutiny faced for their reliance on private data during model training [338, 337]. Additionally, other applications that require model personalization and tuning on private data ranging from spam detection in SMS and emails to word suggestions in keyboards and personal assistant chatbots or HAR, can benefit from edge learning, fostering a paradigm shift towards more ethical and privacy-conscious ML applications.
VII-B4 Reducing energy consumption and carbon footprint
The recent surge in distributed learning methodologies has raised concerns regarding their environmental impact. The substantial energy requirements for training models and data transfer to/from centralized data centers contribute to a significant carbon footprint [412, 413]. Recent trends in ML underscore the critical need to estimate and minimize the environmental impact from the training processes. According to [414], the estimated carbon footprint associated with edge devices by 2027 will be between 22 and 562 MtCO2-eq/year. Therefore, as a pressing research direction, there is a growing emphasis on developing techniques for edge learning that consider energy efficiency. Pioneering works, such as [415], have initiated analyses to quantify the environmental impact of ML in edge devices. Notably, frameworks like FedERA [369], designed for training FL models at the edge, incorporate a dedicated carbon emission tracker module to precisely estimate CO2 emissions during the local parameter update phase.
VII-B5 Frameworks to implement training
@@
Despite the proliferation of frameworks and libraries aimed at enabling ML on edge devices, the current landscape lacks robust support for on-device training. Existing tools, such as PyTorch Mobile and ExecuTorch, predominantly emphasize inference on mobile/edge devices, neglecting the essential backpropagation algorithms crucial for the training phase. This imbalance in focus between training and inference highlights a critical gap in the current ecosystem. Although some tools have been proposed to facilitate on-device training on the edge (see Section VI), there is a pressing need for the development of new libraries, frameworks, and tools explicitly designed for edge learning.
VIII Conclusion
This survey aims to provide a comprehensive overview of the vast field of edge learning, which involves training and fine-tuning of ML models at the edge. We have defined edge learning and its associated metrics and requirements, and explored various techniques and methodologies for optimizing ML training at the edge. Additionally, we explored the growing integration of ML types such as unsupervised learning, reinforcement learning, etc. and the different applications and use cases of edge learning. We have also examined the tools, libraries and frameworks used for edge learning. Furthermore, we have identified key challenges in edge learning and attempted to predict future trends and research directions.
Our analysis has revealed that distributed learning methods, including Federated Learning (FL), are gaining popularity for edge training. We have assessed the benefits and drawbacks of various techniques used to optimize the training at the edge and presented them in Section III-E2 and Table II. We concluded that distributed techniques such as FL and split learning shows great potential for democratizing edge learning. On the other hand, adaptive and fine-tuning based technique should be considered when possible as they often greatly improve the performances or reduce the training time on the edge significantly. Furthermore, model compression techniques are a great choice when slight decreases in performances are acceptable for a reduced model size and computational requirements. At last, we identified a growing trend towards combining different techniques to mitigate their limitations and maximize their benefits.
This survey has provided a broad understanding of edge learning, its requirements, challenges, use cases, and trends, as well as an overview of principal optimization techniques and tools. While it does not provide an in-depth evaluation and comparison of specific task performances, it serves as a reference for developing a foundational understanding of edge learning and identifying areas for future research.